# Enhancing Robustness of CLIP to Common Corruptions through Bimodal Test-Time Adaptation

### v1

@@ 5.2 Ablation Study and Analysis
Adaptation for multiple iterations. We adapt CLIP at test-time with one update per batch. Here, we analyze adaptation for multiple iterations on a single batch. In Figure 5, we illustrate the mean accuracy, across all 15 tasks, using the ViT-B/16 backbone. Continuous adaptation of LayerNorm parameters, which normalize along the features of a single sample, can lead to overfitting as it fits to the noise patterns of the batch, ultimately degrading generalization to other domains. This could also lead to a decline in the loss of CLIP pre-trained knowledge. In the Supplementary, we report the post-adaptation results back on the source test sets - CIFAR10 and CIFAR100. Our approach still maintains a significant advantage over zero-shot CLIP evaluation. Despite a gradual decline in mean accuracy, the performance gap remains notably better, particularly on CIFAR-10C and ImageNet-C. This suggests that the proposed loss components enhance robustness, maintaining effectiveness under challenging conditions.
@@
Contribution of each loss component. In Table 4, we validate the effectiveness of the loss components by taking their different combinations. The mean accuracies across all the tasks are reported using ViT-B/16. The task-wise results are reported in the Supplementary. We observe that the addition of $\mathcal{L}_{pm}$ and $\mathcal{L}_{sp}$ largely improve model performance than entropy minimization via $\mathcal{L}_{ent}$. In fact, the addition of $\mathcal{L}_{sp}$ brings larger improvements, proving that $f_{vis}$ produces discriminative features. Interestingly, using just $\mathcal{L}_{ent}$ achieves better or comparable accuracy than zero-shot CLIP, as the encoders work in synergy while adapting to the domain.
Impact of bimodal adaptation. To show the effectiveness of bimodal adaptation, we ablate $\mathcal{L}_{pm}$ in Eq. 6, which is responsible for updating the text encoder $f_{txt}$. We report the mean classification results using the ViT-B/16 vision backbone in Table 5. As can be seen, we see a drop in accuracy when $f_{txt}$ is “frozen” i.e., when $\mathcal{L}_{pm}$ isn’t used, necessitating the need for bimodal adaptation of CLIP encoders.
Dataset $f_{vis}$ update $f_{txt}$ update Ours
CIFAR-10C 72.58
73.85
CIFAR-100C 41.00
42.09
ImageNet-C 29.88
30.72
@@
Figure 6: Ablation on low batch size - CIFAR-10C and CIFAR-100C (left) and ImageNet-C (right) using ViT-B/16 backbone.
Low batch size setting. In Eq 3, the prototype of each class, via pseudo-labels of each test batch, can be affected by the batch size, with a lower size indicating the absence of a few classes. In Figure 6, we illustrate the mean accuracies on the benchmark datasets. As seen, BAT-CLIP still outperforms or performs comparably to TPT and VTE in low batch size settings. Since TTA happens at the batch-level of a specific domain, the prototypes are computed with the available pseudo-labels, aiding in the continual learning of the feature space as new batches come by.
Effect of different prompt templates. In Table 6, we show results with “relevant” prompt templates to show the independence of such prompt selection, at test-time. As seen, the performance gain over zero-shot ViT-B/32 is fairly large for all the prompt templates. Though TPT fine-tunes a pre-trained prompt on each image, and VTE uses an ensemble of prompts, our method is agnostic to the prompt template being used, making it favorable for real-time usage.
Prompt Template CIFAR-10C CIFAR-100C ImageNet-C
“a low contrast photo of a ¡CLS¿.”
68.53 (+7.81) 37.09 (+4.97) 27.31 (+3.71)
“a blurry photo of a ¡CLS¿.”
68.84 (+10.96) 36.80 (+5.33) 26.92 (+3.52)
“a photo of a big ¡CLS¿.”
67.49 (+10.10) 35.79 (+4.87) 25.64 (+3.29)
@@ 6 Conclusion
In this paper, we propose BAT-CLIP, a bimodal online test-time adaptation framework for CLIP aimed at handling diverse image corruptions simulating real-time environments [15]. Our in-depth analysis of CLIP’s zero-shot performance under increasing corruption severity reveals significant shortcomings in generalization, highlighting the need for effective adaptation. While prior works on TTA for CLIP have predominantly been unimodal focusing on prompt-tuning [38] or prompt ensemble [9] with no model updates, our approach encourages learning rich class-separated visual features via vision encoder updates and strengthens alignment between the image class prototype and corresponding text feature via the text encoder updates. Empirical studies, including ablation results, demonstrate the robustness and substantial performance improvements of our approach over existing methods on benchmark corruption datasets — CIFAR-10C, CIFAR-100C, and ImageNet-C, through synergistic CLIP encoder updates.
@@
In this work, we study the problem of test-time adaptation (TTA) of CLIP [31] towards common image corruptions and propose improved schemes for increasing the robustness of CLIP. We put forward a bimodal domain adaptation scheme, wherein we exploit the shared feature space of CLIP. In essence, leaning towards a more effective multimodal learning and adaptation method, we propose loss components that improve alignment between the class-specific visual prototype and corresponding text features via maximizing the projection. We also increase the cosine distance between the class prototypes to enhance discrimination between visual features. In this Supplementary, we provide additional insights and experimental results, that has been organized as follows,
@@
    Subsection 6.1 offers a detailed discussion of the datasets used, supplemented with visual illustrations.
@@
    To ensure full transparency, we outline the implementation details of all methods in section 6.2, including those for prior TTA approaches [46, 47, 34, 28] adapted for use with CLIP.
@@
    • Subsection 6.3.2 expands on task-specific accuracies, by adopting TTA approaches, for CIFAR-10C, CIFAR-100C, and ImageNet-C, along with a detailed performance discussion.
    • In subsections 6.3.3 and 6.3.4, we present a detailed loss ablation study and the post-adaptation zero-shot generalization on source test sets, respectively — essentially evaluating catastrophic forgetting of BAT-CLIP.
    • Lastly, we include task-wise t-SNE visualizations in subsection 6.3.5 for CIFAR-10C and CIFAR-100C, comparing our method against zero-shot CLIP (ViT-B/16), to illustrate the effectiveness of BAT-CLIP.
@@ 6.1 Datasets
Method
@@
47.98 53.60 60.56 74.30 47.56 73.15 76.43 77.91 79.88 75.66 86.79 71.62 58.34 62.03 64.71 67.37
Ours
ViT-B/32
35.47 39.94 43.23 69.95 41.43 64.50 70.13 70.85 72.33 66.66 81.37 64.57 59.69 48.28 56.62 59.00
TENT
20.09 23.45 34.47 69.85 23.01 39.79 60.35 76.83 77.49 76.07 88.88 81.38 65.35 57.01 51.19 56.35
RoTTA
36.55 40.91 43.99 70.03 42.45 64.52 70.08 71.23 72.68 67.31 81.92 64.99 60.33 49.40 57.11 59.57
RPL
15.89 19.08 34.04 77.84 18.72 41.22 62.39 78.17 78.86 76.31 88.83 81.15 68.98 54.19 51.91 56.51
SAR
50.28 54.12 49.65 73.08 51.98 71.17 74.65 73.73 75.22 70.99 84.25 72.08 63.93 51.57 60.32 65.13
@@
Table 7: Mean accuracy (%) on CIFAR-10C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32. We contrast our results against zero-shot backbones, TENT [46], RoTTA [47], RPL [34], and SAR [28].
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Mean
@@
25.30 27.19 32.78 47.12 23.42 47.16 51.70 51.94 52.48 48.77 61.54 44.50 32.26 33.67 38.06 41.19
Ours
ViT-B/32
16.23 17.83 17.57 39.07 17.63 38.55 43.81 42.32 43.46 39.71 50.32 29.34 28.74 22.85 29.42 31.79
TENT
5.53 7.64 6.85 49.60 4.47 48.45 52.35 49.77 26.77 37.50 63.05 50.53 13.89 27.00 30.80 31.61
RoTTA
16.63 18.25 17.78 38.62 17.76 38.38 43.52 42.39 43.41 39.37 50.60 28.85 28.89 23.50 29.65 31.84
RPL
4.50 5.80 9.61 50.26 4.43 48.88 52.61 50.27 22.36 25.34 63.36 50.31 9.10 18.65 34.53 30.00
SAR
24.63 27.14 21.25 44.57 22.98 43.95 48.40 48.01 47.76 44.85 57.76 42.11 32.69 28.02 33.08 37.81
@@
Table 8: Mean accuracy (%) on CIFAR-100C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32. We contrast our results against zero-shot backbones, TENT [46], RoTTA [47], RPL [34], and SAR [28].
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Mean
@@
17.96 20.46 20.68 25.72 23.04 29.52 26.04 34.92 32.74 39.00 55.00 27.14 19.64 36.66 37.50 29.73
@@
ViT-B/32
12.88 13.04 12.90 24.42 11.86 22.72 20.20 25.70 25.84 30.28 50.54 17.32 18.96 32.20 29.12 23.20
TENT
9.18 8.50 10.42 26.02 15.72 26.06 21.64 27.12 26.18 31.60 50.58 22.28 20.12 34.06 31.30 24.05
RoTTA
13.10 13.30 13.02 24.48 11.96 22.86 20.28 26.06 26.06 30.24 50.46 17.34 19.16 32.44 29.18 23.33
RPL
11.68 10.98 12.10 25.68 13.24 23.98 20.84 26.32 26.12 30.86 50.62 19.30 19.48 33.14 29.92 23.62
SAR
19.82 20.36 20.92 25.78 20.40 28.34 23.10 28.12 28.38 34.74 51.10 24.60 24.38 36.54 34.40 28.07
Ours
16.84 18.20 16.10 25.04 20.90 28.90 25.24 29.42 27.18 36.02 50.18 17.66 27.58 36.20 35.42 27.39
Table 9: Mean accuracy (%) on ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32. We contrast our results against zero-shot backbones, TENT [46], RoTTA [47], RPL [34], and SAR [28].
6.2 Implementation Details
In this section, we summarize the implementation details of all the baseline methods that have been mentioned in the main paper, including ours. We build our approach on the open-source code base  [^3] https://github.com/mariodoebler/test-time-adaptation/tree/main. a standard TTA benchmark codebase, that also houses the hyperparameters and training details of all the prior TTA methods. CLIP-like models are used as provided by OpenCLIP. Only the vision encoder is updated for online TTA methods [46, 47, 34, 28] adopted for CLIP.
BAT-CLIP (Ours): For domain-specific test adaptation, we conducted experiments using ViT-B/16 and ViT-B/32 [10] as the vision backbones. For CIFAR-10C, both the vision encoder ($f_{vis}$) and text encoder ($f_{txt}$) were updated using the AdamW optimizer with a learning rate of 10-3. Similarly, for CIFAR-100C and ImageNet-C, we employed the Adam optimizer and AdamW optimizer, respectively, with a learning rate of 5$\times$10-4. The batch size $\mathcal{B}$ used was set to 200 for CIFAR-10C and CIFAR-100C, and 64 for ImageNet-C. Throughout, the prompt template is fixed to “a photo of a ¡CLS¿.”.
TPT [38]: For each test image, 63 augmentations are generated based on random resized crops, yielding a batch of 64 images, in addition to the original test image. The prompt/context vectors are initialized based on “a photo of a ¡CLS¿.” and tokenized using pre-trained CLIP weights. The confidence threshold is set to 10% i.e., the marginal entropy over the 10% confident samples is minimized. For all the datasets, we follow their core implementation and optimize the prompt vectors using an AdamW optimizer with a learning rate of 5$\times$10-3.
VTE [9]: In VTE, an ensemble of different prompt templates is considered based on the idea of [31]. An example of templates includes “a photo of a ¡CLS¿.”, “a sketch of a ¡CLS¿.”, “a painting of a ¡CLS¿.”, etc. The prompt templates are then averaged. On the vision side, similar to TPT [38], a batch of random augmentation is created for a test image with no model updates.
6.3 Additional Results
6.3.1 Zero-shot performance analysis of ViT-B/32 and ViT-L/14
In the main paper, we analyze and evaluate the zero-shot performance of ResNet-101 (RN101) [14] and ViT-B-16 [10] and conclude that such CLIP backbones are extremely sensitive, in terms of classification accuracy, to increasing severity levels of image corruption. This could be a major concern in situations involving real-time deployment of CLIP. Here, we present a similar analysis in Figure 9, using ViT-B/32 and ViT-L/14 as backbones. Our analysis, from the main paper, carries forward. To summarise, irrespective of the CLIP visual backbone, the robustness towards image corruption is limited. The classification performance degrades with an increase in the severity of corruption in an image.
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Zero-Shot
ViT-B/16
90.1
Ours
ViT-B/32
88.3
Ours
Table 10: Zero-shot performance on CIFAR10 (source) after adaptation of BAT-CLIP on a task.
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Zero-Shot
ViT-B/16
66.6
Ours
ViT-B/32
62.3
Ours
Table 11: Zero-shot performance on CIFAR100 (source) after adaptation of BAT-CLIP on a task.
6.3.2 Results against prior TTA methods
We adopt prior TTA methods [46, 47, 34, 28] for CLIP and present the fine-grained results i.e., the task-wise mean accuracy to demonstrate the efficacy of our method (the higher the better). The results are reported in Tables 7, 8, and 9. Across all the backbones and datasets, we see that our method BAT-CLIP, achieves the best or comparable performance against all the baseline TTA approaches adopted for CLIP.

### v2

@@ 5.3 Ablation Study and Analysis
Adaptation for multiple iterations. Here, we analyze adaptation for multiple iterations on a single batch, shown in Figure 6, using the ViT-B/16 backbone. Continuous adaptation of LN parameters, which normalize along the features of a single sample, can lead to overfitting as it fits to the noise patterns of the batch, ultimately degrading generalization to other domains. This could also lead to a decline in the loss of CLIP pre-trained knowledge. In the Supplementary, we report the post-adaptation results back on the source test sets - CIFAR10 and CIFAR100 to evaluate catastrophic forgetting. Our approach outperforms zero-shot CLIP, with a smaller decline in mean accuracy—especially on CIFAR-10C and ImageNet-C—showing that our loss components improve robustness under challenging conditions.
@@
Ablation on loss components. In Table 3, we ablate the loss components by taking their different combinations using ViT-B/16. The task-wise results are reported in the Supplementary. We observe that the addition of $\mathcal{L}_{pm}$ and $\mathcal{L}_{sp}$ largely improve model performance than entropy minimization via $\mathcal{L}_{ent}$. In fact, the addition of $\mathcal{L}_{sp}$ brings larger improvements, proving that $f_{vis}$ produces discriminative features. Interestingly, using just $\mathcal{L}_{ent}$ achieves better or comparable accuracy than zero-shot CLIP, as the encoders work in synergy while adapting to the domain.
@@
Figure 7: Ablation on low batch size - CIFAR-10C and CIFAR-100C (left) and ImageNet-C (right) using a ViT-B/16 backbone.
Low batch size setting. In Eq 3, the prototype of each class can be affected by the batch size, with a lower size indicating the absence of a few classes. In Figure 7, we illustrate the mean accuracies on the benchmark datasets. As seen, BATCLIP still outperforms or performs comparably to TPT and VTE in low batch size settings. Since TTA happens at the batch-level of a specific domain, the prototypes are computed with the available pseudo-labels, aiding in the continual learning of the feature space as new batches come.
Dataset ZS TPT VTE Ours
CIFAR-10C 75.84 77.75 75.32 84.74
CIFAR-100C 47.82 49.62 48.53 49.25
ImageNet-C 39.55 40.96 40.55 42.84
(a) Mean acc. w/. a ViT-L/14.
Dataset TENT RPL RoTTA SAR Ours
ImageNet-C 40.62 40.00 39.76 42.10 42.84
(b) TTA baselines vs ours w/ ViT-L/14.
Table 4: Ablation on a larger backbone (ViT-L/14) - Mean accuracy (in %)
Extension to larger backbones. We run BATCLIP for a larger backbone like ViT-L/14 and report the results in Table 4. The key advantage lies in the projection matching loss ($\mathcal{L}_{pm}$) which updates the text encoder to produce more image-aware text features—an aspect missing in other state-of-the-art TTA baselines.
Computational efficiency. Due to an additional text encoder update, BATCLIP updates $\sim 0.044$% of all CLIP parameters, with a ViT-B/16 backbone, as opposed to $\sim 0.026$% in other TTA baselines. Though the increase is marginal, this leads to superior performances as adaptation of the text encoder produces input-aware CLIP text features, aligning better with the image features. On an NVIDIA RTX A5000 GPU, it takes 0.2s/batch on ImageNet-C with a batch size of 64, compared to 2.34s for WATT [36]. Also, WATT adapts per template (for several iterations), making it expensive every time. BATCLIP achieves faster on-the-fly adaptation rates and is favorable for real-time deployment.
@@ 6 Conclusion
In this work, we propose BATCLIP, a bimodal online test-time adaptation framework for CLIP aimed at handling diverse image corruptions simulating real-time environments. Our in-depth analysis of CLIP’s zero-shot performance under increasing corruption severity reveals significant shortcomings in generalization, highlighting the need for effective adaptation. Although prior work on TTA for CLIP has predominantly been unimodal focusing on prompt-tuning or prompt ensemble with no model updates, BATCLIP, through synergistic CLIP encoder updates, encourages learning rich class-separated visual features via vision encoder updates and strengthens alignment between the image class prototype and corresponding text feature via text encoder updates. Empirical studies, including ablations, show robust performance gains over all baselines on benchmark corruption and domain generalization datasets.
@@
In this work, we study the problem of online test-time adaptation (TTA) of CLIP [38] towards common image corruptions and propose improved schemes for increasing the robustness of CLIP. In addition, to demonstrate the broad impact of our proposed approach, we evaluate on common domain generalization datasets [14]-OfficeHome [51], PACS [25], VLCS [12], and Terra Incognita [3]. We put forward a bimodal domain adaptation scheme, at test-time, wherein we exploit the shared feature space of CLIP. In essence, leaning towards a more effective multimodal learning and adaptation method, we propose loss components that improve alignment between the class-specific visual prototype and corresponding text features via maximizing the projection. We also increase the cosine distance between the class prototypes to enhance discrimination between visual features. In this Supplementary, we provide additional insights and experimental results, that has been organized as follows,
@@
    Section 6.1 offers a detailed discussion of the standard common corruption datasets [18] used, supplemented with visual illustrations.
@@
    To ensure full transparency, we outline the implementation details of all the methods in Section 6.2, including those for prior TTA approaches [53, 54, 41, 35] adapted for CLIP. We also discuss details of the adapted version of WATT [36] - WATT-P* and WATT-S* for online TTA, and other details of experiments run on the domain generalization datasets.
@@
    • In Subsection 6.3.2, we report the main online TTA results on CIFAR-10C, CIFAR-100C, and ImageNet-C with a ViT-B/32 backbone.
    • We study the effect of different prompt templates on BATCLIP in Subsection 6.3.4. Subsection 6.3.5 discusses the ablation of $\mathcal{L}_{pm}$ which is responsible for updating the text encoder.
    • In subsections 6.3.3 and 6.3.6, we present a detailed loss ablation study and the post-adaptation zero-shot generalization on source test sets, respectively — essentially evaluating catastrophic forgetting of BATCLIP.
    • Lastly, we include task-wise t-SNE visualizations in subsection 6.3.7 for CIFAR-10C and CIFAR-100C, comparing our method against zero-shot CLIP (ViT-B/16), to illustrate the effectiveness of BATCLIP.
@@ 6.1 Datasets
6.2 Implementation Details
In this section, we summarize the implementation details of all the baseline methods that have been mentioned in the main paper, including ours. We build our approach on the standard benchmark code base  [^2] https://github.com/mariodoebler/test-time-adaptation/tree/main that also houses the hyperparameters and training details of all the prior TTA methods. CLIP-like models are used as provided by OpenCLIP. Only the vision encoder is updated for existing online TTA methods [53, 54, 41, 35] adopted for CLIP.
6.2.1 Experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C
BATCLIP (Ours): For domain-specific test adaptation, we conducted experiments using ViT-B/16 and ViT-B/32 [11] as the vision backbones. For CIFAR-10C, both the vision encoder ($f_{vis}$) and text encoder ($f_{txt}$) were updated using the AdamW optimizer with a learning rate of 10-3. Similarly, for CIFAR-100C and ImageNet-C, we employed the Adam optimizer and AdamW optimizer, respectively, with a learning rate of 5$\times$10-4. The batch size $\mathcal{B}$ used was set to 200 for CIFAR-10C and CIFAR-100C, and 64 for ImageNet-C. Throughout, the prompt template is fixed to “a photo of a $<$CLS$>$.”.
TPT [45]: For each test image, 63 augmentations are generated based on random resized crops, yielding a batch of 64 images, in addition to the original test image. The prompt/context vectors are initialized based on “a photo of a $<$CLS$>$.” and tokenized using pre-trained CLIP weights. The confidence threshold is set to 10% i.e., the marginal entropy over the 10% confident samples is minimized. For all the datasets, we follow their core implementation and optimize the prompt vectors using an AdamW optimizer with a learning rate of 5$\times$10-3.
VTE [10]: In VTE, an ensemble of different prompt templates is considered based on the idea of [38]. An example of templates includes “a photo of a $<$CLS$>$.”, “a sketch of a $<$CLS$>$.”, “a painting of a $<$CLS$>$.”, etc. The prompt templates are then averaged. On the vision side, similar to TPT [45], a batch of random augmentation is created for a test image with no model updates.
WATT-P* and WATT-S* [36]: The two original variants of WATT [36] were proposed with weight-averaging of adapted weights from multiple prompt templates. Additionally, for each test batch, adaptation was performed over multiple iterations. To fit our online TTA scheme, we reduced the number of iterations to a single step for each prompt template and reset CLIP parameters only after a domain. However, this is still not fully online, as training is performed on the test batch using 8 selected prompt templates: “a photo of a $<$CLS$>$”, “itap of a $<$CLS$>$”, “a bad photo of the $<$CLS$>$”, “a origami $<$CLS$>$”, “a photo of the large $<$CLS$>$”, “a $<$CLS$>$ in a video game”, “art of the $<$CLS$>$”, and “a photo of the small $<$CLS$>$”. As they report performance on CIFAR-10C and CIFAR-100C only, we follow their original implementation details. For experiments using WATT-S*, we set a batch size of 200 (for a fair comparison to other baselines) and a learning rate of 10-3 using an Adam optimizer. For WATT-P*, a learning rate of 10-4 is used with the same batch size.
StatA [55]: We adopt the original hyperparameter settings of StatA for online TTA in our reported ImageNet experiments and apply them to ImageNet-C. To control the effective number of correlated classes in a test batch, we set $\gamma$ to 0.1 (low correlation) and -1 (separate, sequential). The default prompt template is “a photo of a $<$CLS$>$.”
6.2.2 Experiments on Domain Generalization Datasets
We evaluate BATCLIP on standard domain generalization datasets [14]. For a fair comparison, as reported in WATT [36], we use a batch size of 128 for all the online TTA experiments on VLCS, PACS, and Office Home. We use an AdamW optimizer for model updates using BATCLIP with learning rates of 5$\times$10-4, 10-3, and 5$\times$10-3 for OfficeHome, PACS, and VLCS, respectively. We use the same learning settings for TENT, WATT-S*, and WATT-P*, as in WATT [36]. In addition to the mentioned datasets, we also run experiments on the Terra Incognita dataset and optimize using an AdamW optimizer with a learning rate of 5$\times$10-4 and a batch size of 256.
6.3 Additional Results
6.3.1 Zero-shot performance analysis of ViT-B/32 and ViT-L/14
In the main paper, we analyze and evaluate the zero-shot performance of ResNet-101 (RN101) [17] and ViT-B/16 [11] and conclude that such CLIP backbones are extremely sensitive, in terms of classification accuracy, to increasing severity levels of image corruption. This could be a major concern in situations involving real-time deployment of CLIP. Here, we present a similar analysis in Figure 10, using ViT-B/32 and ViT-L/14 as backbones. Our analysis, from the main paper, carries forward. To summarise, irrespective of the CLIP visual backbone, the robustness towards image corruption is limited. The classification performance degrades with an increase in the severity of corruption in an image.
Method Venue
@@ 6.3.1 Zero-shot performance analysis of ViT-B/32 and ViT-L/14
ICML’22 50.28 54.12 49.65 73.08 51.98 71.17 74.65 73.73 75.22 70.99 84.25 72.08 63.93 51.57 60.32 65.13
[…]
TPT
NeurIPS’22 43.11 46.53 48.29 71.31 47.80 66.89 71.96 74.00 76.00 68.81 84.12 66.35 63.86 51.86 58.01 62.59
VTE
ECCV-W’24 47.59 50.18 53.15 71.39 53.86 67.92 72.90 76.37 76.30 70.78 83.27 61.07 69.00 58.57 61.14 64.90
WATT-P*
NeurIPS’24 43.64 47.1 45.97 74.98 48.04 70.42 74.74 76.1 76.86 71.85 85.15 70.36 64.6 56.0 60.37 64.41
WATT-S*
NeurIPS’24 56.38 58.08 52.48 78.07 58.29 76.42 79.16 78.9 79.71 76.76 87.19 76.49 70.35 64.11 65.05 70.49
@@
ICML’22 24.63 27.14 21.25 44.57 22.98 43.95 48.40 48.01 47.76 44.85 57.76 42.11 32.69 28.02 33.08 37.81
[…]
TPT
NeurIPS’22 16.08 17.65 17.54 39.21 19.47 38.91 44.01 43.45 44.46 40.15 50.93 27.77 30.91 23.36 29.55 32.23
VTE
ECCV-W’24 16.84 18.33 18.94 39.63 22.88 39.13 43.80 44.56 44.88 39.21 49.37 28.37 34.13 26.87 30.12 33.14
WATT-P*
NeurIPS’24 15.55 17.02 16.16 40.25 16.16 38.74 43.6 42.49 43.51 39.39 51.17 31.85 28.35 23.94 29.74 31.86
WATT-S*
NeurIPS’24 16.22 17.72 16.85 41.54 17.04 39.66 44.55 43.33 44.26 40.26 52.13 33.13 29.34 24.65 30.39 32.73
@@
ICML’22 19.82 20.36 20.92 25.78 20.40 28.34 23.10 28.12 28.38 34.74 51.10 24.60 24.38 36.54 34.40 28.07
[…]
TPT
NeurIPS’22 12.04 12.64 12.52 25.38 12.28 22.68 20.78 26.36 26.64 30.78 51.02 16.50 19.90 33.62 30.62 23.58
VTE
ECCV-W’24 11.96 12.32 13.44 25.06 11.70 22.58 22.40 27.38 27.02 32.28 51.52 16.84 19.94 34.80 32.82 24.14
StatA ($\gamma$=0.1)
CVPR’25 11.57 12.28 11.74 21.91 10.91 20.43 18.87 24.04 25.0 29.12 49.21 16.02 18.87 29.15 26.72 21.73
StatA ($\gamma$=-1)
CVPR’25 12.52 13.10 12.74 22.43 11.39 21.17 19.61 24.37 25.60 29.44 49.89 17.71 19.21 29.87 27.52 22.44
@@
Table 5: Mean accuracy (%) on CIFAR-10C, CIFAR-100C, and ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/32.
6.3.2 Online TTA results using a ViT-B/32 backbone
In the main paper, we had presented the online TTA results on CIFAR-10C, CIFAR-100C, and ImageNet-C, using a ViT-B/16 backbone. In Table 5, we provide results using a ViT-B/32 backbone. Across all the datasets, we see that BATCLIP achieves the best or comparable performance against all the baseline approaches.